CRY+VIS Interface Meeting
Date: 2022/1/14 13:00-13:30
Participants: Takahashi, Yokozawa, Ikeda, Ushiba, Aso, Uchiyama, Sato, Hirata
- IX: Health check. Spectrum was improved. Investigation of the glitch in the F3 signals is still ongoing. Windshield for the MN oplev is not working well.
- EX: Health check.
EY: ALIGN mode is available. LT->Yaw coupling is large. Preparation for ACCs.
- Replacement of flip cables: SRM has not been finished yet. Cables for Type-C will be delivered in the end of Jan.
- TM height setting: ITMY.
- Oplev: Setup of the TM Oplev in IY after the TM alignment.
- ACC: Preparation for EY.
- Commissioning in air: Return of GRY. Alignment of SR3/ETMY.
- Commissioning in vacuum: Go to the stable lock.
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Cabling of the power line in IXV/EXC on 1/11-12.
- Heater control: Test in IXV/EXV is planed. Repaired power controller will be delivered on 1/17.
- Power source for Oplev:
- Tunnel for the IYC booth: Planned on 1/31.
Date: 2022/1/14 14:00-
Participants: Takahashi, Ushiba, Fabian, Ishizaki, Aso, Uchiyama, Terrence
- Preparing for the ACCs in EY. The tuning of the resonant frequency was finished.
- Fixed the cables around the IP in EY.
- [Terrence] We aligned the EY and get the reflected green under the following condition:
- IP L, T, controlled.
- BF Y, controlled.
- Coarse alignment of yaw using BF yaw pico motor.
- Coarse alignment of pitch using MN pitch actuation.
- There's something strange. I can't engage all IP controls at the same time. It used to be okay. I will check them later.
- Terrence reported SR3 wasn't healthy, the TF measurements looked bad.
- We inspected the suspension and found several security structure screws touching the IM or IRM.
- This suspension was healthy, how did it become unhealthy?
- As they should be, the screws were likley close to the IM and IRM and began touching the payload when yaw was adjusted with the F0 stepper motor.
- When we removed the F0 yaw stepper motor we followed a method to keep the same yaw position, but it seems it was not accurate enough.
- We lost the alignment give by the oplev.
- The incident angle of the green beam might have changed also.
- Terrence updated several control filters of the SR3 and modified guardian.
- In the beginning, I thought the yaw oscillation was due to misbehaving control loops at the IP so I modified them in the beginning but turned out it was not helping.
- I modified the rest of the control filters for consistency.
- They inspected the PR3 traverser to check whether it had the locking assenmblies or they had been removed.
- We calibrated PRM SF LVDT. The calibration factor hasn't been written in the medm screen yet.
- In PR2, we connected the cable for TM-H4. A cable with the same label has beed connected. There are old cables around that flange and they should be removed.
- Type-A team: Preparation for the ACCs in EY, Heter test in IXV/EXV, Check of the control in EY, Height adjustment of ETMY.
- Type-B team: BS SF calibration, Y-arm commissioning, SRM flip cable
(Date in Kamioka)
The week from 1/17
- Takahashi: 1/18-21
- Sato: 1/18-21
- Hirata: 1/18-21
- Aso: 1/17-20
The week from 1/24
- Takahashi: 1/24-28
- Sato: 1/24-28
- Hirata: 1/24-28